OpenStudy (anonymous):

find the derivatives of the given functions 5. g(x)=x^4+3x^3-8x-7

OpenStudy (anonymous):

repeated applications of the power rule \(\frac{d}{dx}x^n=nx^{n-1}\)

OpenStudy (anonymous):

do you mean the answer is 8x-3

OpenStudy (jdoe0001):

no, see your exponents, as follow as satellite73 suggested

OpenStudy (jdoe0001):

you just have a straight polynomial, so the power rule by itself will do

OpenStudy (anonymous):

so what is ur answer

OpenStudy (jdoe0001):

\(\huge ax^n = (n\times a)x^{n-1}\)

OpenStudy (jdoe0001):

let's try the 1st term, you have \(\bf x^4\\ ax^n = (n\times a)x^{n-1}\\ 1x^4 = (1 \times 4)x^{4-1}\)

OpenStudy (anonymous):

oh ok. i had 4x^3+9x^2 - 8

OpenStudy (anonymous):

im i right?

OpenStudy (jdoe0001):

\(8x\\ ax^n = (n\times a)x^{n-1}\\ 8x^1 = (1 \times 8)x^{1-1}\\ x^0 = 1\\ 8\times 1 = 8 \) yes :)

OpenStudy (jdoe0001):

and 7 is a constant and the derivative of a constant is 0

OpenStudy (anonymous):

oh ok.fanks so much